A rectangular room is 1.2 times as long as it is wide, and its perimeter is
30 meters. Find the dimension of the room.

Answer by ikleyn(52778)

.

x + 1.2x + x + 1.2x = 30


4.4x = 30


x = 30/4.4 = 6.818 meters.


ANSWER.  The width is  6.82 m;  the length is  1.2*6.82 = 8.18 m.